OpenStudy (countrygirl1431):

x - y = 6 x = y + 2 Please Help!!! Solve the system of equations by substitution. (4, -2) (6, 0) no solution coincident

OpenStudy (alexandervonhumboldt2):

x=6+y x=y+2 6+y=y+2 0=4 impossible thus no solution

OpenStudy (anonymous):

First, put them in the same form (variables in same order)

OpenStudy (countrygirl1431):

Thanks @AlexandervonHumboldt2

OpenStudy (anonymous):

no solution because when you put (4,-2) it is right but when you put it in the second one it is wrong